OpenStudy (anonymous):

What is the product of the product of 2 times the square root of 35 x and 3 times the square root of 14 x? Simplify if possible. @radar :)

OpenStudy (anonymous):

\[2\sqrt{35x} \times 3\sqrt{14x}\]

OpenStudy (anonymous):

answer choices: \[294x \sqrt{10}\] \[42x \sqrt{10}\] \[12x^2\sqrt{10}\] \[42\sqrt{10x^2}\]

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Well let's break it apart to make it a bit easier, can you do that?

OpenStudy (anonymous):

Expand it? o.o

OpenStudy (luigi0210):

Yea, like break the numbers inside the sqrt's. Like if we broke this up it would be: